OpenStudy (anonymous):

write an equation in slope intercept form (if possible) for each line through (1,-1), m=-2

OpenStudy (anonymous):

Y=mx+c

OpenStudy (anonymous):

so y=-2x

OpenStudy (anonymous):

Yea I think so

OpenStudy (anonymous):

not quite correct...

OpenStudy (anonymous):

you have the slope part right... m = -2, so the x-term is correct as "-2x"

OpenStudy (anonymous):

but y = -2x goes through (0,0), (1,-2), (2,-4), (3,-6), etc.... not (1,-1)

OpenStudy (anonymous):

when you have a slope and a point, you use the well-named "point-slope" form of a line equation: if the point is (x1, y1) and the slope is m, then the point-slope form is y - y1 = m(x - x1) So you substitute in the point (x1, y1) in the places above where you see x1 and y1, and you put in your slope m (in this problem, it is = -2) for the "m" so your point is (1, -1), making x1 = 1 and y1 = -1 and slope is m = -2: y - (-1) = -2(x - 1) now just simplify...

OpenStudy (anonymous):

@teresaa I gave you some good help on your point slope line equation problem... sorry I missed you before you went offline.... good luck!!